Domino Co. has the following data related to an item of inventory: Beginning Inventory, March 1 - 100 units @ $2.10 Purchase, March 7 - 350 @ $2.20 Purchase, March 16 - 70 @ $2.25 Ending Inventory, March 31 - 130

The value assigned to cost of goods sold if Domino uses FIFO is

The value assigned to ending inventory if Domino uses LIFO is
Answer #1
Units sold 390 =100+350+70-130
Value assigned to cost of goods under FIFO 848 =(100*2.10)+(390-100)*2.20
Value assigned to ending inventory under LIFO 276 =(100*2.10)+(130-100)*2.20